Buffy is rolling along in her 13.0 kg wagon at 3.0 m/s (in the positive direction) when she jumps off the back. She continues to move forward at 1.5 m/s relative to the ground. This causes her wagon to go speeding forward at 7.00 m/s relative to the ground. How much does Buffy weigh?
Answer #1

let mass of buffy be M.

then using conservation of linear momentum:

(13+M)*3=M*1.5+13*7

M=34.667 kg
Answer #2

SOLUTION :


Let Buffy weighs W kg .


So, momentum of both Buffy and wagon before separation

= (W + 13) * 3 

= 3 W + 39   kg-m/s 


Momentum after Buffy separates from the wagon 

= W * 1.5 + 13 * 7 

= 1.5 W + 91 kg-m/sec


Momentum remains unchanged, so,


3 W + 39 = 1.5 W + 91

=> (3 - 1.5) W = 91 - 39 

=> 1.5 W = 52

=> W = 52/1.5 

=> W = 34.67 kg


So, weight of Buffy is 34.67 kg (ANSWER).

answered by: Tulsiram Garg